# ReGAL: Refactoring Programs to Discover Generalizable Abstractions

### v1

@@ 7 Discussion and Conclusion
Connections to Semantic Parsing.   Executable semantic parsing (Winograd, 1972; Zelle & Mooney, 1996) typically involves mapping queries to a domain-specific language (DSL); the DSL specifies a set of abstractions that are useful for a particular goal, e.g. SQL operations for querying databases, or spatial abstractions for robotics. These DSLs are typically defined by a human; one way to view ReGAL is as a way of learning a DSL on top of an extremely general set of primitives. One of the benefits of ReGAL is its generality: on three different domains, it is able to learn useful abstractions without human intervention, while, in a standard semantic parsing setting, these abstractions would need to be manually specified.
@@
Connections to Hierarchical Reinforcement Learning.   One way to view the functions discovered by ReGAL is as low-level policies – or skills – composed of primitive actions. In this view, our system is similar to hierarchical reinforcement learning (HRL; Barto & Mahadevan, 2003), where tasks are factorized into skill selection and skills themselves. In hierarchical frameworks, there is typically a high-level controller policy that selects lower-level policies. In our setting, the agent LLM acts as a controller, selecting from a set of skills, while ReGAL’s training stage is responsible for discovering a useful set of skills; this is akin to option discovery (Sutton et al., 1999), where closed-loop policies for specific skills are learned from a reward signal. While ReGAL has a similar hierarchical structure, it differs from HRL in that ReGAL’s skills are symbolic, interpretable, and editable, as opposed to HRL skill policies, which typically have none of these features.
@@
Limitations. As mentioned in connection to HRL, the functions ReGAL learns are code-based. This can make them less flexible than functions parameterized by neural networks (e.g. Andreas et al. (2016)), especially in domains where the environment can change dynamically. For instance, consider a navigation domain where an agent needs to get to the kitchen in various homes; depending on the primitive actions, a function that succeeds in one home will likely fail in another. However, ReGAL’s verification-based pruning means that no function would be discovered for this navigation task. Relatedly, not every domain has reusable abstractions, and not every example stands to benefit from them. In many cases, the primitives for a domain are the right level of abstraction for that domain, e.g. if they already form a DSL. The abstractions are also not always ideal; in Sec. B.1 we see that ReGAL’s abstractions are not necessarily the same as those a human would choose, e.g. a human would likely write a function like $\tt{draw\_5gon(size)}$ rather than $\tt{draw\_small\_5gon\text{()}}$).
@@
Conclusion. We introduce ReGAL, a gradient-free approach to learning abstractions from a small set of examples. Our experimental results show that abstractions from ReGAL improve the accuracy of programs predicted by a variety of LLMs across three diverse domains. Furthermore, ReGAL abstractions are reusable and general, allowing them to be applied across examples for a given task. In our analysis, we find that the functions learned by ReGAL codify commonly-used subroutines as well as task dynamics. Our error analysis indicates that ReGAL’s improvements come both from function reuse as well as simplification of the reasoning involved in program prediction.
8 Broader Impacts
@@
Acknowledgements
We thank Yichen Jiang, Justin Chen, Jaehong Yoon, and Swarnadeep Saha for their valuable feedback on the paper. This work was supported by NSF-AI Engage Institute DRL-2112635, DARPA Machine Commonsense (MCS) Grant N66001-19-2-4031, and Accelerate Foundation Models Research program. The views contained in this article are those of the authors and not of the funding agency.
@@ A.3 Testing
In our test-time agent, we use ChromaDB for indexing and retrieval [^6] https://github.com/chroma-core/chroma/ with OpenAI Ada embeddings. ICL examples are retrieved from the training data and from the Demo Bank using query similarity. We limit the number of Code Bank functions to 20, using the similarity between the query and the function name for retrieval. The Code Bank is pruned once before testing.
@@ A.5 Data
Dataset Train Dev Test # Primitives
LOGO 200 100 111 9
Date 66 113 180 4
TextCraft 190 50 77 3
Table 4: Dataset statistics. We list the number of primitive functions in the programs (aside from built-in Python functions).
@@ A.5.1 LOGO
LOGO data is generated from a synchronous text-code grammar, and pairs procedurally-generated language commands like “three small triangles in a row” with a corresponding Turtle graphics program; however, the original LOGO dataset is expressed in a Lisp-style functional syntax. While this facilitates the application of helpful data structures for efficient code search (Ellis et al., 2021; Bowers et al., 2023), object-oriented languages like Python are far more common in practice. As a result, they are represented more in LLM pretraining data, which has been shown to contribute to parsing performance (Bogin et al., 2023). To account for this, we write an AST-based parsing script to translate the LOGO dataset into Python.
@@
$\tt{forward(dist)}$
move forward $\tt{dist}$ units
$\tt{left(theta)}$
rotate left by $\tt{theta}$ degrees
$\tt{right(theta)}$
rotate right by $\tt{theta}$ degrees
$\tt{penup\text{()}}$)
lift the pen (stop drawing)
$\tt{pendown\text{()}}$)
put the pen down (start drawing)
$\tt{teleport(x,y,theta)}$
move to position $\tt{(x,y)}$ with angle $\tt{theta}$
$\tt{heading\text{()}}$)
get the current angle of the turtle
$\tt{isdown\text{()}}$)
check if the pen is down
$\tt{embed(program,vars)}$
runs the code in $\tt{program}$ using the current context $\tt{vars}$ and teleports back to the original position.
Table 5: LOGO Primitives
@@ A.5.2 Date understanding
Date understanding involves both mathematical reasoning and parsing. Each question poses a word problem that requires reasoning about dates and times. For example, problems ask questions like: “On May 9th, 2017 Jane bought 40 eggs. She ate one per day. Today she ran out of eggs. What is the date 10 days ago in MM/DD/YYYY?”. These kinds of questions are especially hard for LLMs to answer directly. Lyu et al. (2023) approached this task as a program prediction task, wherein an LLM predicts a Python script that gives the answer to the question when executed. This paradigm is especially helpful for Date, as there are existing Python libraries that can perform math on dates, such as $\tt{datetime}$ and $\tt{dateutil}$. While predicting programs with these libraries results in strong performance as compared to LLM-based reasoning, Lyu et al. (2023) method predicts programs one at a time, leaving the benefits of shared sub-routines on the table. We use the programs predicted by Lyu et al. (2023) as a starting point for our refactoring process. Table 6 describes the Python libraries called by Lyu et al. (2023)’s programs, which we treat as the primitives for Date.
@@
$\tt{date\text{()}}$)
returns a $\tt{date}$ object
$\tt{time\text{()}}$)
returns a $\tt{time}$ object
$\tt{relativedelta(time)}$
performs addition/subtraction of $\tt{time}$, which can be days, weeks, months, or years.
$\tt{strftime(format)}$
prints the date in the specified format
Table 6: Date Primitives
@@ A.5.3 TextCraft
The environment has three primitives: $\tt{inventory}$, $\tt{craft}$, and $\tt{get}$ which we convert into Python variants, described in Table 7.
@@
$\tt{getObject(obj\_name)}$
get $\tt{obj\_name}$ from the environment
$\tt{craftObject(obj\_name,}$ $\tt{\hskip 30.00005pt[ingredients])}$
craft $\tt{obj\_name}$ using the list of ingredients
$\tt{checkInventory\text{()}}$)
return the contents of the inventory
Table 7: TextCraft Primitives
@@ B.1 What kinds of programs are discovered
Table 8, Table 9, and Table 10 show examples of the discovered programs most commonly used by the agent.
    def draw_small_5gon():
@@
Table 8: Examples of discovered programs for LOGO
    def get_date_today(date_obj):
@@
Table 9: Examples of common discovered programs for Date
    def craft_object_with_ingredients(target,
@@
Table 10: Examples of common discovered programs for TextCraft
@@ Appendix C Hyperparameters
ICL budget split 0.5 0.5 0.5
Table 11: Hyperparameter settings for all experiments
@@ Appendix D Prompts
Please rewrite the following two programs to be more efficient.
{primitive description string}
The resulting programs MUST execute to the same result as the original programs.
Start by writing helper functions that can reduce the size of the code.
You can also choose from the following helper functions:
{codebank function definitions}
// repeated for all in batch
QUERY {i}: {query}
PROGRAM {i}: {program}
Please format your answer as:
// repeated for i
NEW PROGRAM {i}:
// once at end
NEW HELPERS:
Do not include any text that is not valid Python code.
Recall that no matter what, your program MUST be formatted in the following fashion:
// repeated for i
NEW PROGRAM {i}:
# Thoughts:
# 1. The query asks for: <query intention>
2. <query> can be solved by <components>.
# 3. I will use helper function <function> to <goal>.
<code for program {i}>
Try to make your new programs as short as possible by introducing shared helper functions. Helper function parameters should be as general as possible and helper functions should be informatively named.
{logo_special_instr}
@@
You are an expert coder. For each query below, decompose it into its parts.
Example:
Query: Do some action 5 times and then do another action
Query (decomposed):
The query asks: Do some action and then do another action
This can be decomposed into:
1. repeat an action
2. some action
3. another action
Query: {query}
Query (decomposed):
@@
Please add comments to the following program to explain what each chunk of code does with respect to the query.
First, decompose the query into parts. Then comment the code with the query parts.
Example:
Query: Do some action and then do another action
Code:
do_some_action()
do_another_action()
Query: Do some action 5 times and then do another action
Query (decomposed):
The query asks: Do some action and then do another action
This can be decomposed into:
1. repeat an action
2. some action
3. another action
Commented code:
# repeat an action
for i in range(5):
    # do some action
    do_some_action()
# do another action
do_another_action()
{primitive description}
Query: {query}
Code:
{program}
Query (decomposed):
{output from decompose prompt}
@@
Refactor the following function to improve performance.
FUNCTION:
{func_str}
{library_str}
You may also use the following helper functions:
{codebank_str}
Try to increase the number of passing programs. Try to make programs general. For example, you can add parameters instead of hardcoded values or call other helper functions. First, for each failing query, explain why the programs do not accomplish the querys goal. Output this reasoning as:
Thoughts:
1. The function passes some tests and fails others because <reason>.
2. The failing queries <repeat queries here> asked for <intent>.
3. The program failed because <reason>.
4. This can be addressed by <change>.
Then output your program so that all test cases pass, using the following format: NEW PROGRAM: <program>
@@
SUCCEEDED:
{example of passing case}
FAILED:
{example of failing case}
Thoughts:
@@
Your task is to solve simple word problems by creating Python programs.
{codebank_str}
You will be given a query and have to produce a program. {thought_str}
Examples:
{icl_string}
Please generate ONLY the code to produce the answer and nothing else.
Query: {query}
{thought_and}Program:
@@
Your task is to draw simple figures using python Turtle graphics.
You will use a custom turtle library, similar to the built-in library, which is sufficient for all tasks.
Heres a description of the custom library:
- forward(x): move forward x pixels
- left(theta): rotate left by theta degrees
- right(theta): rotate right by theta degrees
- penup(): stop drawing
- pendown(): start drawing
- teleport(x, y, theta): move to position (x, y) with angle theta
- heading(): get the current angle of the turtle
- isdown(): check if the pen is down
- embed(program, local_vars): runs the code in program using the current context and teleports back to the original position. Allows you to nest programs. Implementationally, embed gets the turtle state (is_down, x, y, heading), executes program, then returns to the original state.
- save(path): save the picture to file
{codebank_str}
@@
Examples:
{icl_string}
Please generate ONLY the code to produce the answer and nothing else.
Query: {query}
Thought and Program:

### v2

@@ 7 Discussion
Connections to Semantic Parsing.   Executable semantic parsing (Winograd, 1972; Zelle & Mooney, 1996) typically involves mapping queries to a domain-specific language (DSL), a set of abstractions for a particular application, e.g., SQL operations for querying databases. These DSLs are manually defined; one way to view ReGAL is as a way of learning a DSL on top of an extremely general set of primitives. A key benefit of ReGAL is its generality: on five different domains, it learns useful abstractions without human intervention, while, in a standard semantic parsing setting, these abstractions would designed by hand.
@@
Connections to Hierarchical Reinforcement Learning.   Another way to view the functions discovered by ReGAL is as low-level policies composed of primitive actions. In this view, ReGAL resembles hierarchical reinforcement learning (HRL; Barto & Mahadevan, 2003), where tasks are split into skill selection by a controller and the skill policies themselves. Our agent LLM acts as a controller while ReGAL’s training stage is responsible for discovering a useful set of skills; this is akin to option discovery (Sutton et al., 1999). While ReGAL has a similar hierarchy, it differs in that its skills are symbolic, interpretable, and editable, as opposed to HRL policies, which typically are not.
@@
Limitations. As mentioned in connection to HRL, the functions ReGAL learns are code-based. This can make them less flexible than functions parameterized by neural networks (e.g., Andreas et al., 2016), especially in domains where the environment can change dynamically, e.g., navigation tasks. However, ReGAL’s verification-based pruning means that no functions would be discovered in these cases. Relatedly, not every domain has reusable abstractions, and not every example stands to benefit from them; the primitives for a domain may already be suitably abstract, e.g., if they already form a DSL. Finally, in Sec. B.1 we see that ReGAL’s abstractions are not necessarily the same as those a human would choose.
8 Conclusion.
@@
We introduce ReGAL, a gradient-free approach to learning abstractions from a small set of examples. Our experimental results show that abstractions from ReGAL improve the accuracy of programs predicted by a variety of LLMs across five diverse domains. Furthermore, ReGAL abstractions are reusable and general, allowing them to be applied across examples for a given task. In our analysis, we find that the functions learned by ReGAL codify commonly-used subroutines as well as task dynamics. Our error analysis indicates that ReGAL’s improvements come both from function reuse as well as simplification of the reasoning involved in program prediction.
Acknowledgements
We thank Yichen Jiang, Justin Chen, Jaehong Yoon, and Swarnadeep Saha, as well as the anonymous reviewers, for their valuable feedback on the paper. This work was supported by NSF-AI Engage Institute DRL-2112635, DARPA Machine Commonsense (MCS) Grant N66001-19-2-4031, and the Accelerate Foundation Models Research program. The views contained in this article are those of the authors and not of the funding agencies.
Impact Statement
@@ A.2 Training
Table 5: Dataset statistics. We list the number of primitive operations in the programs (aside from built-in Python functions).
Dataset Train Dev Test # Primitives
LOGO 200 100 111 9
Date 66 113 180 4
TextCraft 190 50 77 3
MATH (Alg.) 194 61 74 0
TabMWP 194 60 74 0
@@ A.3 Testing
In our test-time agent, we use ChromaDB [^6] https://github.com/chroma-core/chroma/ for indexing and retrieval with OpenAI Ada embeddings. ICL examples are retrieved from the training data and from the Demo Bank using query similarity. We limit the number of Code Bank functions to 20, using the similarity between the query and the function name for retrieval. The Code Bank is pruned once before testing.
@@ A.5.1 LOGO
LOGO data is generated from a synchronous text-code grammar, and pairs procedurally-generated language commands like “three small triangles in a row” with a corresponding Turtle graphics program; however, the original LOGO dataset is expressed in a Lisp-style functional syntax. While this facilitates the application of helpful data structures for efficient code search (Ellis et al., 2021; Bowers et al., 2023), object-oriented languages like Python are far more common in practice. As a result, they are represented more in LLM pretraining data, which has been shown to contribute to parsing performance (Bogin et al., 2023). To account for this, we write an AST-based parsing script to translate the LOGO dataset into Python. The LOGO dataset was released under an MIT License.
@@
$\tt{forward(dist)}$ move forward $\tt{dist}$ units
$\tt{left(theta)}$ rotate left by $\tt{theta}$ degrees
$\tt{right(theta)}$ rotate right by $\tt{theta}$ degrees
$\tt{penup\text{()}}$) lift the pen (stop drawing)
$\tt{pendown\text{()}}$) put the pen down (start drawing)
$\tt{teleport(x,y,theta)}$ move to position $\tt{(x,y)}$ with angle $\tt{theta}$
$\tt{heading\text{()}}$) get the current angle of the turtle
$\tt{isdown\text{()}}$) check if the pen is down
$\tt{embed(program,vars)}$ runs the code in $\tt{program}$ using the current context $\tt{vars}$ and teleports back to the original position.
@@ A.5.2 Date understanding
Date understanding involves both mathematical reasoning and parsing. Each question poses a word problem that requires reasoning about dates and times. For example, problems ask questions like: “On May 9th, 2017 Jane bought 40 eggs. She ate one per day. Today she ran out of eggs. What is the date 10 days ago in MM/DD/YYYY?”. These types of questions are especially hard for LLMs to answer directly. Lyu et al. (2023) approached this task as a program prediction task, wherein an LLM predicts a Python script that gives the answer to the question when executed. This paradigm is especially helpful for Date, as there are existing Python libraries that can perform math on dates, such as $\tt{datetime}$() and $\tt{dateutil}$(). While predicting programs with these libraries results in strong performance as compared to LLM-based reasoning, Lyu et al. (2023) method predicts programs one at a time, leaving the benefits of shared sub-routines on the table. We use the programs predicted by Lyu et al. (2023) as a starting point for our refactoring process. Table 7 describes the Python libraries called by Lyu et al. (2023)’s programs, which we treat as the primitives for Date. Date was released under an Apache 2.0 License.
Table 7: Date Primitives
@@
$\tt{date\text{()}}$) returns a $\tt{date}$ object
$\tt{time\text{()}}$) returns a $\tt{time}$ object
$\tt{relativedelta(time)}$ performs addition/subtraction of $\tt{time}$, which can be days, weeks, months, or years.
$\tt{strftime(format)}$ prints the date in the specified format
@@ A.5.3 TextCraft
The environment has three primitives: $\tt{inventory}$, $\tt{craft}$, and $\tt{get}$ which we convert into Python variants (Table 8). This dataset uses the Apache 2.0 license. To obtain primitive programs in the train set, we use all the depth-2 instances not in the test set, and use the ADaPT (Prasad et al., 2023) trajectories with $d=4$. We then perform rule-based translation of environment commands into calls to primitive operations and filter out erroneous actions (based on the saved textual environment feedback) and all thought statements.
Table 8: TextCraft Primitives
@@
$\tt{getObject(obj\_name)}$ get $\tt{obj\_name}$ from the environment
$\tt{craftObject(obj\_name,}$ $\tt{\hskip 30.00005pt[ingredients])}$ craft $\tt{obj\_name}$ using the list of ingredients
$\tt{checkInventory\text{()}}$) return the contents of the inventory
A.5.4 MATH
MATH (Hendrycks et al., 2021) contains challenging math word problems with open-ended answers (i.e. not multiple-choice). Unlike LOGO, Date, and TextCraft, MATH does not have any primitives, as each problem can be solved using standard Python math tools (addition, subtraction, etc.); this makes MATH a challenging setting for discovering helper functions, as the space of programs is less constrained.
A.5.5 TabMWP
Like MATH, TabMWP (Lu et al., 2022) is a dataset of word problems. In this case, the word problems pertain to tabular data, where each problem asks questions involving computation over tabular data. This dataset has been used in prior work on tool learning, e.g. Yuan et al. (2023). Like MATH, there is no existing domain-specific language for TabMWP.
@@ B.1 What kinds of programs are discovered
Figs. 5, 6 and 7 show examples of the discovered programs most commonly used by the agent.
@@
Figure 5: Examples of discovered programs for LOGO as mentioned in Figs. 1 and 3. As the name suggests, $\tt{draw\_small\_5gon}$() draws a small-size pentagon and $\tt{draw\_semicircle}$() draws a small semicircle.
@@
Figure 6: Examples of common discovered programs for Date as mentioned in Fig. 3. All the helpers functions and variables are named intuitively reflection their functionality.
@@
Figure 7: Examples of common discovered programs for TextCraft as shown in Figs. 1 and 2. $\tt{craft\_object\_with\_ingredients}$() encapsulate the game dynamics of TextCraft as it first fetches the inventory, and every ingredient not in the inventory is obtained from the environment prior to using the crafting the target object. Similarly, the helper $\tt{check\_and\_get\_object}$() gets an object from the environment if it is not already in the inventory.
x = (point1[0] + point2[0]) / 2
y = (point1[1] + point2[1]) / 2
return (x, y)
def calculate_square(num):
return num ** 2
Figure 8: Examples of discovered programs for MATH. $\tt{calculate\_midpoint}$ is an example of reuse, where ReGAL finds a frequently-used functionality and encapsulates it, while in $\tt{calculate\_square}$ ReGAL wraps a fairly easy function with a more informative name.
total_cost = 0
for item in items:
total_cost += packages[item]
return total_cost
def add_prices(prices):
total_cost = 0
for price in prices.values():
total_cost += price
return total_cost
Figure 9: Examples of discovered programs for TabMWP, reflecting the domain, which often involves costs and prices.
@@ Appendix C Hyperparameters
Figure 10: Validation performance of CodeLlama-13B with ReGAL abstractions trained using different batch sizes.
C.1 Varying Batch Size
Fig. 10 shows the performance of a CodeLlama-13B agent on the Date dev set for ReGAL abstractions trained using different batch sizes. We see that performance can vary substantially and non-linearly with batch size, with 3 being the best setting and 5 a close second.
Table 9: Hyperparameter settings for all experiments
@@ C.1 Varying Batch Size
ICL budget ratio 0.5 0.5 0.5
@@ Appendix D Prompts
Please rewrite the following two programs to be more efficient. {primitive description string} The resulting programs MUST execute to the same result as the original programs. Start by writing helper functions that can reduce the size of the code. You can also choose from the following helper functions: {codebank function definitions} // repeated for all in batch QUERY {i}: {query} PROGRAM {i}: {program} Please format your answer as: // repeated for i NEW PROGRAM {i}: // once at end NEW HELPERS: Do not include any text that is not valid Python code. Recall that no matter what, your program MUST be formatted in the following fashion: // repeated for i NEW PROGRAM {i}: # Thoughts: # 1. The query asks for: <query intention> 2. <query> can be solved by <components>. # 3. I will use helper function <function> to <goal>. <code for program {i}> Try to make your new programs as short as possible by introducing shared helper functions. Helper function parameters should be as general as possible and helper functions should be informatively named. {logo_special_instr}
@@
You are an expert coder. For each query below, decompose it into its parts. Example: Query: Do some action 5 times and then do another action Query (decomposed): The query asks: Do some action and then do another action This can be decomposed into: 1. repeat an action 2. some action 3. another action Query: {query} Query (decomposed):
@@
Please add comments to the following program to explain what each chunk of code does with respect to the query. First, decompose the query into parts. Then comment the code with the query parts. Example: Query: Do some action and then do another action Code: do_some_action() do_another_action() Query: Do some action 5 times and then do another action Query (decomposed): The query asks: Do some action and then do another action This can be decomposed into: 1. repeat an action 2. some action 3. another action Commented code: # repeat an action for i in range(5): # do some action do_some_action() # do another action do_another_action() {primitive description} Query: {query} Code: {program} Query (decomposed): {output from decompose prompt}
@@
Your task is to solve simple word problems by creating Python programs. {codebank_str} You will be given a query and have to produce a program. {thought_str} Examples: {icl_string} Please generate ONLY the code to produce the answer and nothing else. Query: {query} {thought_and}Program: